# Cooperative and Asynchronous Transformer-based Mission Planning for Heterogeneous Teams of Mobile Robots

### v1

@@ IV-C Asynchronous Multi-Agent Transformer
IV-C1 Macro-Observations Embedder
@@
During the centralized training phase, this module transforms a sequence of macro-observations $(z^{i_{1}},\dots,z^{i_{m}})$ into a sequence of macro-observation embeddings $(z_{e}^{i_{1}},\dots,z_{e}^{i_{m}})$ to be used as input tokens for the encoder. In the distributed execution phase, each agent $i_{l}\in\mathcal{N}(i)$ inputs its macro-observation $z^{i_{l}}$ into the trained macro-observations embedder on its own hardware, then transmits the resulting embedding $z_{e}^{i_{l}}\in\mathbb{R}^{d}$ to its neighborhood’s broker. Transmitting MO embeddings, vectors of length $d$, instead of actual MOs, which in our case are local and global multi-channeled maps of dimensions $L\times L\times 4$ and $S\times S\times 8$, significantly reduces network traffic. The broker subsequently passes the full sequence of embeddings $(z_{e}^{i_{1}},\dots,z_{e}^{i_{m}})$ to the encoder.
The architecture of the macro-observations embedder is shown in Fig. 3, where two separate convolutional neural networks (CNNs) extract features from the multi-channeled global and local maps, concatenate and feed them into a multi-layer perceptron (MLP). An adaptive max pooling layer is used after the first CNN to scale the global feature maps of any $S\times S$ dimensions to $G\times G$ zones, allowing the model to work in any environment size. To enable the learning of unique behaviors for different classes of agents, the output of the MLP is combined via element-wise addition with a learnable class encoding obtained by passing a one-hot class identifier of size $C$ through a fully connected (FC) layer with output size $d$.
@@
Figure 3: The macro-observation embedder network, where agent $i$’s macro-observation $z^{i}$ is transformed into a macro-observation embedding $z_{e}^{i}\in\mathbb{R}^{d}$
IV-C2 Encoder
@@
The encoder is made up of several encoding blocks each consisting of a self-attention mechanism, an MLP, and residual connections. The sequence of MO embeddings $z_{e}^{i_{1:m}}$ are fed as the input to these computational blocks, where they are encoded into a sequence of MO representations $(\hat{z}^{i_{1}},\dots,\hat{z}^{i_{m}})$. These latent representations carry information both about each agent’s current view of the environment, as well as the high-level interrelationships among the agents. An additional MLP is also used during the training phase to approximate the value of each agent’s observation. Values associated with the active subset of agents, $V(\hat{z}^{i_{1}}),\dots,V(\hat{z}^{i_{g}})$ are used to minimize the empirical Bellman error
$L_{Encoder}(\phi)=\frac{1}{T}\sum_{\tau=0}^{T-1}\frac{1}{g(\tau)}\sum_{l=1}^{g(\tau)}\Bigl{[}R_{\tau}^{i_{l}}+\gamma V_{\bar{\phi}}(\hat{z}_{\tau+1}^{i_{l}})\\ -V_{\phi}(\hat{z}_{\tau}^{i_{l}})\Bigr{]},$ (3)
@@
where $T$ is the total number of MA updates, $g(\tau)$ is the number of active agents at the $\tau$th MA update instance, and $\bar{\phi}$ is the non-differentiable target network’s parameter.
IV-C3 Macro-Actions Embedder
@@
This module consists of an MLP and an fully connected layer. The MLP converts a one-hot encoded representation of MAs $m^{i_{0:l-1}},l=\{1,\dots,m\}$ into MA embeddings $m_{e}^{i_{0:l-1}}$, which form a sequence of vectors with dimension $d$. Similar to the MO embedder, the FC layer generates learnable class encodings, which are combined with the MA embeddings to associate each macro-action with the agent class responsible for executing it.
IV-C4 Decoder
@@
The decoder processes the joint MA embeddings $m_{e}^{i_{0:l-1}}$, $l=\{1,\dots,m\}$ through a series of decoding blocks, with $m_{e}^{i_{0}}$ acting as an arbitrary token designating the start of decoding. Each decoding block is made up of a masked self-attention mechanism, a masked attention mechanism, and an MLP followed by residual connections. The masking ensures that each agent is only attending to itself and the agents preceding it, preserving the sequential updating scheme and the monotonic performance improvement guaranty during training [10]. The final decoder block outputs a sequence of joint MA representations $\{\hat{m}^{i_{0:h-1}}\}_{h=1}^{l}$, which is then fed to an MLP to obtain the probability distribution of agent $i_{l}$’s MA, which is the high-level policy $\mu^{i_{l}}_{\theta}(m^{i_{l}}|\hat{z}^{i_{1:m}},\hat{m}^{i_{0:l-1}})$, where $\theta$ represents the decoder parameters. The decoder is trained by minimizing the following clipped PPO objective, which only uses the action probabilities and advantage estimates of the active subset of agents $i_{1:g(\tau)}$ at the $\tau$th instance of MA updates.
$L_{Decoder}(\theta)=-\frac{1}{T}\sum_{\tau=0}^{T-1}\frac{1}{g(\tau)}\sum_{l=1}^{g(\tau)}\min(r_{\tau}^{i_{l}}(\theta)\hat{A}_{\tau}^{i_{l}},\\ clip(r_{\tau}^{i_{l}}(\theta),1\pm\epsilon)\hat{A}_{\tau}^{i_{l}}),$ (4)
@@
IV-D High-Level Processor
During centralized training, the high-level processor module manages agent activation, forms macro-observations, and updates the navigation controller with the current global goal. In the execution phase, this module handles communication and the exchange of MO embeddings and MAs with the neighbors as well.
IV-E Low-Level Navigation Controller
This module consists of a global path planner and a motion controller that generates the robot’s primitive action, $a^{i}_{t}$, at each time step $t$. A path planning algorithm, such as D* Lite [38], could be employed to determine the shortest path from the robot’s current location to the global goal specified by its MA $m^{i}$. This path is represented as a sequence of coordinates on the global occupancy grid map. The motion controller may use a local planner [39] to optimize a local trajectory based on the global path and the robot’s sensory data, ensuring the trajectory meets the robot’s motion constraints, avoids obstacles, and minimizes execution time. The optimized trajectory is then converted into low-level motion control commands, forming the primitive action $a^{i}_{t}$.
IV-F Collaborative Simultaneous Localization and Mapping
This module enables robots to collaboratively build a shared global occupancy grid map. A state-of-the-art algorithm like Swarm-SLAM [40] can be utilized during execution, where agents may experience intermittent communication. A designated robot temporarily processes data from connected robots to generate the most accurate map and pose estimates, ensuring convergence to a unified global reference frame after multiple rendezvous among robot subsets. The discretized occupancy grid map is then employed by the high-level processor to construct macro-observations and by the path planner to set navigation goals.
V Simulation Setup
V-A Environment, Agents and Rewards
@@
CATMiP is evaluated in a customized 2D grid-world environment designed in Minigrid [41], with randomly generated environments for each episode, featuring varied numbers of cluttered rooms with different shapes, and a randomly positioned static target. The environment provides clues to the target’s whereabouts in the form of detectable traces in nearby cells.
@@
Fig. 4 shows a rescuer robot (red triangle) and two explorer robots (blue and yellow triangles) searching for the target (green square). Each robot’s global goal is marked by a colored circle. Cells are either free space (black) or occupied by walls or objects (orange circles), with explored areas highlighted. Free spaces that contain clues are colored pink. Robots move at different speeds: explorer robots traverse a cell in one time step, while the rescuer robot takes two time steps for actions like moving forward or turning.
@@
Figure 4: A snapshot of a simulated episode in the Minigrid environment, where two explorer agents and one rescuer agent navigate an unknown area in search of a target.
@@
The main objective is for a rescuer robot to reach the target as quickly as possible. Each agent $i\in I$ receives a time-dependent reward $r^{i}_{success}(t)=300(1-0.9\frac{t}{h})$ upon the mission’s completion, incentivizing strategies that result in a quick rescue. Additionally, a similar time-dependent team reward $r^{i}_{locate}(t)=100(1-0.9\frac{t}{h})$ is given at the time step the target is discovered. Class-specific intrinsic rewards are introduced to enhance the learning process in an environment with such sparse rewards. Before the target is located, an exploration reward $r^{i}_{explore}(t)=\frac{50}{n_{free}}n^{i}_{t}$ is given to the explorer agent $i$, where $n_{free}$ represents the total number of free cells in the map and $n^{i}_{t}$ is the number of cells discovered solely by agent $i$ at time $t$. Once the target is found, the rescuer agent $i$ incurs a small penalty $r^{i}_{distance}(t)=-1$ at each time step $t$ that it does not move closer to the target.
V-B Macro-Action and Macro-Observation Spaces
@@
Global goal candidates, or macro-actions, are defined as cells within a square of side length $L$ centered on the agent, giving rise to a discrete macro-action space of size $L^{2}$. Once a macro-action is selected, it is mapped to its corresponding coordinate $(x,y)$ on the grid map, establishing the agent’s global goal. Invalid actions, such as those targeting occupied or out-of-bounds cells, are handled using invalid action masking [42]. This has proven effective in large action spaces such as those in StarCraft II [43] and Dota 2 [44].
@@
Each agent’s macro-observation is composed of three elements: a global information map of size $S\times S\times 8$, a local information map of size $L\times L\times 4$, and a one-hot encoded agent class identifier. The channels in the local map provide information about cells in the agent’s immediate vicinity, including whether they have been explored, their occupancy, the presence of the target or any clues leading to it, and the agent’s last or current navigation goal. Similarly, the global map’s first four channels represent this information for the entire map, while the remaining four channels encode the agent’s current location, its decaying trajectory, the locations of other rescuer robots, and the locations of other explorer robots.
@@
In Minigrid, the primitive action space includes four actions: moving forward, turning right, turning left, and stopping. An agent’s primitive actions are selected to follow the shortest path to the global goal.
VI Simulation Results
We evaluated the effectiveness of the proposed CATMiP framework using the Minigrid simulation environment under varying environmental conditions and communication constraints. Simulations were performed on three tasks, each featuring different map sizes and team compositions. Task 1 involved one rescuer and one explorer agent in a $16\times 16$ grid. Task 2 increased the map size to $20\times 20$, with an additional randomly selected agent. Task 3 scaled up to a $32\times 32$ environment with 8 agents, consisting of 2 rescuers, 4 explorers, and 2 randomly assigned agents.
Two different models with the same network architecture and training hyperparameters were trained on Tasks 1 and 2, utilizing 128 parallel environments. Training was conducted on an NVIDIA GeForce RTX™ 3090 GPU over 250 million total steps, equivalent to 9765 episodes with a horizon of 200 steps. Model 1 required approximately 52 hours of training, while Model 2 took 96 hours. Fig. 5 shows the progression of the average mission success rate throughout training, where an exponential moving average with a span of 200 is applied to the data. Key hyperparameters were set as follows: $L=7$, $G=4$, $\gamma=1$, with 10 training epochs and a learning rate of $10^{-4}$.
@@ VI Simulation Results
Figure 5: Progression of the mission success rate against the training steps for the two models.
VI-1 Scalability
To assess the scalability of CATMiP with respect to environment size and team composition, we tested the two trained models across all three tasks, conducting 500 simulated episodes per task with identical random seeds and no communication loss. Fig. 6 shows the probability of success as a function of elapsed time since the start of the mission. Both models demonstrated the ability to scale effectively, handling larger teams and environments. Notably, Model 2 outperformed Model 1 across all tasks, particularly in Task 3, which required higher levels of team coordination. This improvement can be attributed to the additional training experiences per episode and the development of more refined coordination strategies through training with multiple agents of each class.
Figure 6: The success rates of Models 1 and 2 on tasks with varying environment and team sizes.
VI-2 Communication
Model 2 was tested across 500 simulated episodes under four communication conditions: no communication loss, $\sigma=6$, $\sigma=4$, and $\sigma=2$. Note that a lower $\sigma$ value corresponds to a steeper decline in communication probability as a function of the distance between agents. As shown in Fig. 7, even under severe communication constraints, where agents lose connection beyond 5 unit distances, they are still able to complete the mission within a reasonable time frame.
Figure 7: Performance of Model 2 is evaluated on Task 3 with increasing levels of communication loss.
VII Conclusions & Future Work

### v2

@@ 5 Asynchronous Multi-Agent Transformer (AMAT)
5.1 Macro-Observations Embedder
@@
During centralized training, this module transforms a sequence of MOs $(z^{i_{1}},\dots,z^{i_{m}})$ into embeddings $(z_{e}^{i_{1}},\dots,z_{e}^{i_{m}})$ as input tokens for the encoder. In the distributed execution phase, each agent $i_{l}\in\mathcal{N}(i)$ locally processes its MO $z^{i_{l}}$ using the trained embedder, then transmits the resulting embedding $z_{e}^{i_{l}}\in\mathbb{R}^{d}$ to its neighborhood’s broker. This approach reduces network traffic by transmitting compact embeddings instead of high-dimensional MOs, which include local and global maps.
The architecture, shown in Fig. 3, uses two separate convolutional neural networks (CNNs) to extract features from multi-channeled global and local maps. An adaptive max pooling layer scales global feature maps of any $S\times S$ dimensions to fixed $G\times G$ zones, enabling the model’s operation in varying environment sizes. The features are then concatenated and processed by a multi-layer perceptron (MLP).
Similar to the idea of agent indication [25], a learnable agent class encoding is used to enable unique behaviors for different classes of agents. This class encoding is obtained by processing a one-hot agent class identifier tied to each agent’s MO $z^{i_{l}}$ through a fully connected layer (FC). The outputs of the MLP and FC layers are combined via element-wise addition, resulting in the macro-observation embedding $z^{i_{l}}_{e}$.
@@
Figure 3: The macro-observation embedder network, where agent $i_{l}$’s macro-observation $z^{i_{l}}$ is transformed into a macro-observation embedding $z_{e}^{i_{l}}\in\mathbb{R}^{d}$
5.2 Encoder
@@
The encoder is made up of several encoding blocks each consisting of a self-attention mechanism, an MLP, and residual connections. It processes the sequence of MO embeddings $z_{e}^{i_{1:m}}$ into a sequence of MO representations $(\hat{z}^{i_{1}},\dots,\hat{z}^{i_{m}})$, which carry information both about each agent’s current view of the environment, as well as the high-level interrelationships among the agents. An additional MLP is also used during the training phase to approximate the value of each agent’s observation. Values associated with the active subset of agents, $(V_{\phi}(\hat{z}^{i_{1}}),\dots,V_{\phi}(\hat{z}^{i_{g}}))$, are used to minimize the empirical Bellman error
$L_{Encoder}(\phi)=\frac{1}{T}\sum_{\tau=0}^{T-1}\frac{1}{g(\tau)}\sum_{l=1}^{g(\tau)}\Bigl{[}R_{\tau}^{i_{l}}+\gamma V_{\bar{\phi}}(\hat{z}_{\tau+1}^{i_{l}})-V_{\phi}(\hat{z}_{\tau}^{i_{l}})\Bigr{]},$ (3)
@@
where $T$ is the total number of MA updates, $g(\tau)$ is the number of active agents at the $\tau$th MA update instance, $\phi$ represents MO-embedder and encoder parameters, and $\bar{\phi}$ represents the non-differentiable target network’s parameters.
5.3 Macro-Actions Embedder
@@
This module converts one-hot encoded representation of MAs $m^{i_{0:l-1}},l=\{1,\dots,m\}$ into MA embeddings $m_{e}^{i_{0:l-1}}$ using an MLP. Similar to the MO embedder, class encodings are combined with these embeddings to associate each MA with the agent class responsible for executing it.
5.4 Decoder
@@
The decoder processes the joint MA embeddings $m_{e}^{i_{0:l-1}}$, $l=\{1,\dots,m\}$ through a series of decoding blocks, with $m_{e}^{i_{0}}$ acting as an arbitrary token designating the start of decoding. Each decoding block is made up of a masked self-attention mechanism, a masked attention mechanism, and an MLP followed by residual connections. The masking ensures that each agent is only attending to itself and the agents preceding it, preserving the sequential updating scheme and the monotonic performance improvement guaranty during training [26]. The final decoder block outputs a sequence of joint MA representations $\{\hat{m}^{i_{0:h-1}}\}_{h=1}^{l}$, which is then fed to an MLP to obtain the probability distribution of agent $i_{l}$’s MA, which is the high-level policy $\mu^{i_{l}}_{\theta}(m^{i_{l}}|\hat{z}^{i_{1:m}},\hat{m}^{i_{0:l-1}})$, where $\theta$ represents the MA-embedder and decoder parameters. The decoder is trained by minimizing the following clipped PPO objective, which only uses the action probabilities and advantage estimates of the active subset of agents $i_{1:g(\tau)}$ at the $\tau$th instance of MA updates.
$L_{Decoder}(\theta)=-\frac{1}{T}\sum_{\tau=0}^{T-1}\frac{1}{g(\tau)}\sum_{l=1}^{g(\tau)}\min(r_{\tau}^{i_{l}}(\theta)\hat{A}_{\tau}^{i_{l}},clip(r_{\tau}^{i_{l}}(\theta),1\pm\epsilon)\hat{A}_{\tau}^{i_{l}}),$ (4)
@@
6 Simulation Setup
6.1 2D Simulation Environment
@@ 6 Simulation Setup
CATMiP is evaluated in a customized 2D grid-world environment built using Minigrid [34]. Each episode features randomly generated environments of size $S\times S$, consisting of cluttered rooms with diverse shapes and a static target placed at an unknown location.
@@
A snapshot of the simulation is shown in Fig. 4, where a rescuer robot (red triangle) and two explorer robots (blue and yellow triangles) navigate the environment to locate the target (green square). Each robot’s global goal is marked by a corresponding colored circle. Cells are either free space (black) or occupied by walls or objects (orange circles), with explored areas visually highlighted. Free cells adjacent to the target are marked pink and serve as detectable clues for the robots. Robots move at different speeds: explorer robots traverse a cell in one time step, while the rescuer robot takes two time steps for actions like moving forward or turning.
@@
Figure 4: A snapshot of a simulated episode in the Minigrid environment, where two explorer agents (yellow and blue triangles) and one rescuer agent (red triangle) navigate an unknown area in search of a target (green square). The pink squares are detectable as clues to the target’s location. Agents’ current navigation goals are marked by corresponding colored circles.
6.2 Reward Structure
@@
The main objective is for a rescuer robot to reach the target as quickly as possible. Each agent $i\in I$ receives a time-dependent reward $r^{i}_{success}(t)=300(1-0.9\frac{t}{h})$ upon the mission’s completion, incentivizing strategies that result in a quick rescue. Additionally, a similar time-dependent team reward $r^{i}_{locate}(t)=100(1-0.9\frac{t}{h})$ is given to all agents at the time step the target is discovered. To conserve energy and reduce unnecessary movement, agents incur a small penalty of $r^{i}_{movement}(t)=-0.05$ whenever they move to a new cell.
6.3 Macro-Action and Macro-Observation Spaces
@@
Global goal candidates, or macro-actions, are defined as cells within a square of side length $L$ centered on the agent, giving rise to a discrete macro-action space of size $L^{2}$. Once a macro-action is selected, it is mapped to its corresponding coordinate $(x,y)$ on the grid map, establishing the agent’s global goal. Invalid actions, such as those targeting occupied or out-of-bounds cells, are handled using invalid action masking [35, 36].
@@
Each agent’s macro-observation is composed of three elements: a global information map of size $S\times S\times 7$, a local information map of size $L\times L\times 6$, and a one-hot encoded agent class identifier. The channels in the local map provide information about cells in the agent’s immediate vicinity. These channels indicate which cells have been explored, the cells’ occupancy, the location of the target and clues around it, the agent’s current navigation goal, the locations of other rescuer robots, and the location of other explorer robots. Similarly, the global map represents this information for the entire map, with an extra channel representing the agent’s current location.
During centralized training or perfect communication conditions, these channels contain the latest information from other agents. However, during distributed execution when there is no communication link between explorer agent $i$ and rescuer agent $j$ for example, current location of agent $j$ would not show up in agent $i$’s rescuer agents’ location channel and agent $i$’s location would not be visible in agent $j$’s explorer agents’ location channel.
@@
In Minigrid, the primitive action space includes four actions: moving forward, turning right, turning left, and stopping. An agent’s primitive actions are selected to follow the shortest path to the agent’s current navigation goal generated by the A* algorithm.
7 Simulation Results
We evaluated the effectiveness of the proposed CATMiP framework in the Minigrid simulation environment under varying environmental conditions, team compositions, and communication constraints. To assess CATMiP’s performance, we compared it against both multi-agent and single-agent planning-based exploration methods adapted for multi-agent scenarios.
7.1 Training Setup
Asynchronous Training: CATMiP was trained using asynchronous centralized training on an NVIDIA GeForce RTX™ 3090 GPU. The training was conducted across 64 parallel environments for 62,500 episodes with an episode horizon of 200 steps. Each episode included 3 agents: one rescuer agent, one explorer agent, and a third agent randomly assigned to either class. Macro-actions had a maximum duration of 10 time steps, after which agents automatically transitioned to standby mode. The asynchronous training process took approximately 285 hours.
Synchronous Training (Synch-CATMiP): For comparison, we trained a synchronous version of CATMiP, where macro-actions were updated for all agents simultaneously every 10 time steps. This synchronous training approach follows a procedure similar to MAT [26]. Due to fewer trajectories being collected and stored in the training buffer, training over the same total number of episodes required 166 hours.
In both cases, key hyperparameters were set as follows: embedding size $d=192$, local decision-making range $L=7$, global pooling range $G=4$, discount factor $\gamma=1$, a linearly decaying learning rate starting from $10^{-4}$, and using 10 training epochs. Both models have 1,280,786 total trainable parameters. Fig. 5 shows the progression of the average mission success rate and the average agent reward throughout the training for both models, where an exponential moving average with a span of 200 was applied for visualization. It can be seen that both models follow a similar trend in training for 62,500 episodes.
@@ 7.1 Training Setup
Figure 5: (a) Progression of the mission success rate and (b) progression of average agent rewards against the number of episodes during training of the two models.
7.2 Baseline Methods
We compare our method with planning-based exploration methods, including the multi-agent method of artificial potential field (APF) [37] and three single-agent frontier-based methods, namely a utility-maximizing algorithm (Utility) [38], a search-based nearest frontier method (Nearest) [39], and a rapid-exploring-random-tree-based method (RRT) [40]. The single-agent methods are adapted to multi-agent settings by planning on the shared global map, and their implementation on the Minigrid environment is adapted from Yu et al. [20]. During communication dropouts, the shared global map contains only the latest information received from other agents. All agent classes were treated identically during exploration. However, once a rescuer agent detects the target’s location, it immediately navigates along the shortest path to the target.
7.3 Evaluation Results and Analysis
The trained CATMiP models (both asynchronous and synchronous) and the planning-based baselines were evaluated on three tasks with increasing complexity. Task 1 involved one rescuer and one explorer agent in a $15\times 15$ grid. Task 2 increased the map size to $20\times 20$, with an additional explorer agent in the team. Task 3 scaled up to a $32\times 32$ environment with 6 agents consisting of 2 rescuers and 4 explorers. Evaluations are performed over 100 different episodes with the same random seed. Agents acted asynchronously during all evaluations.
Figure 6 shows the success rate of the different methods against mission time in Task 1. This comparison is made for three cases with different communication constraints. In Figure 6 (a) the agents have consistent communication throughout the mission, whereas in (b) and (c) the value of $\sigma$ in Equation 2 is set to $4$ and $2$ respectively, indicating increasing levels of communication loss. The learning-based models outperform the baselines in all three cases by showing higher success rates within the same time-frame. For example, as seen in Figure 6 (a), $97\%$ of the experiments using the CATMiP and Synch-CATMiP model were successful by the 100th step since the start of the mission, while the best performing baseline method, RRT, achieves $86\%$ success by the same time. Since the models were trained for a more complex task, this shows the scalibility of CATMiP to smaller environments and team sizes.
The same comparison between the models with different communication constraints is shown for Task 2 in Figure 7. Once again, the learning-based models outperform the baselines in all three communication scenarios.
For Task 3, with a map size of $32\times 32$ and $6$ agents, the learning-based models still show top performance alongside the planning-based methods RRT and Nearest, as shown in Figure 8. As the communication loss increases, CATMiP’s superiority becomes more prominent. Since CATMiP and Synch-CATMiP were trained on a smaller map of size $20\times 20$ and with a smaller team of agents, results on Task 3 show the scalability of our proposed framework to more complex tasks, with larger maps and team sizes. Moreover, in both Tasks 2 and 3, the CATMiP model which was trained with the asynchronous training scheme shows better adaptability to asynchronous execution than Synch-CATMiP, especially as the communication loss increases and agents have to act mostly based on their own observations only.
Figure 6: Success rate of different models on Task 1, with different communication constraints: (a) distributed execution with no communication loss, (b) distributed execution with moderate communication loss, and (c) distributed execution with heavy communication loss.
Figure 7: Success rate of different models on Task 2, with different communication constraints: (a) distributed execution with no communication loss, (b) distributed execution with moderate communication loss, and (c) distributed execution with heavy communication loss.
Figure 8: Success rate of different models on Task 3, with different communication constraints: (a) distributed execution with no communication loss, (b) distributed execution with moderate communication loss, and (c) distributed execution with heavy communication loss.
8 Conclusions & Future Work